1.   The height of grade 1 pupils is approximately normally distributed with µ = 45 inches and s = 2.

a. If an individual pupil is selected at random, what is the probability that he or she has a height of 42 and 47?

b. A class of 30 of these pupils is used as a sample. What is the probability that the class mean is between 42 and 47?

c.  If a pupil is selected at random, what is the probability that is taller than 46 inches?

d. A class of 30 of these pupils is used as sample. What is the probability that the class mean is greater than 46 inches?





1
Expert's answer
2022-03-19T02:40:38-0400

Population mean"~\mu=45"

Standard deviation"~\sigma=2"


(a) "P(42~and~47)"

"Z=\dfrac{X-\mu}{\sigma}"

"Z_{42}=\dfrac{42-45}{2}=-1.5"

"P_{42}=0.06681"

"Z_{47}=\dfrac{47-45}{2}=1.0"


"P_{47}=1-0.84134"


"P_{47}=0.15866"


"=0.06681 ~~and~~0.15866"


(b) "P(42<\bar X<47)"


"=P(\frac{42-45}{2}<Z<\frac{47-45}{2})"


"=P(-1.5<Z<1.0)"


"=P(Z<1.0)-P(Z<-1.5)"


"=0.77453"


(c) "P(X>46)=P(Z>\frac{46-45}{2})"


"=P(Z>0.5)"


"=1-P(Z<0.5)"


"=0.30854"


(d) "P(\bar X>46)=P(Z>\frac{46-45}{2})"


"=P(Z>0.5)"


"=1-P(Z<0.5)"


"=0.30854"